OpenStudy (anonymous):

Circle A has an equation of x2 + (y + 3)2 = 16. Where is the center of the circle located?

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] has center (h,k) radius r

OpenStudy (anonymous):

you have \[(x-0)^2+(y-(-3))^2=4^2\] so center is \[(0,-3)\] radius if 4

OpenStudy (anonymous):

x^2+6y=16